# Relativistic and Electron Correlation Effects in Static Dipole Polarizabilities for Main-Group Elements

### v1

@@ 1 Introduction
A recent study by Schwerdtfeger and Nagle compiled an updated table of the most accurate calculated and experimental dipole polarizabilities for neutral atoms in the Periodic Table, with nuclear charges ranging from $Z=1$ to $120$, except for livermorium (Lv, $Z=116$).3 However, there remains potential for improvement in the polarizabilities of open-shell elements and heavy atoms that exhibit significant relativistic effects. This study focuses on all main-group atoms, except for hydrogen (H, $Z=1$), particularly for heavy atoms (i.e., atoms with $Z>40$).
@@
The determination of accurate atomic polarizabilities requires high-level electron correlation treatment within a relativistic framework, particularly for heavy atoms. Over the past two decades, researchers have advanced techniques to incorporate spin-orbit coupling (SOC) effects in dipole polarizability calculations. For closed-shell atoms and open-shell atoms with only one open-shell electron, the coupled cluster theory with single and double substitutions (CCSD) with perturbative triples, CCSD(T), is usually employed within a fully relativistic Dirac-Coulomb framework, where both scalar-relativistic effects and SOC effects can be included. Due to the SOC effect, atoms in Group 14 are also treated as closed-shell atoms. However, for open-shell atoms with more than one open-shell electron in Groups 15–17, the coupled cluster theory is not yet well-developed at the fully relativistic level, and an alternative is multi-reference configuration interaction (MRCI).3
@@
Previous studies have thoroughly investigated the polarizabilities considering SOC effects at the CCSD or CCSD(T) level for cesium (Cs, $Z=55$)4, 5 and francium (Fr, $Z=87$)4, 6, 7 in Group 1. Fleig et al. examined the spin-orbit-resolved static polarizabilities of Group 13 atoms, including gallium (Ga, $Z=31$), indium (In, $Z=49$), and thallium (Tl, $Z=81$), using four-component configuration interaction (CI) and coupled cluster methods with ANO-RCC basis sets in the Dirac-Coulomb framework.8 Borschevsky et al. later improved these results using the Fock space coupled cluster (FSCC) with more extensive Fægri basis sets.4 Pershina et al. first reported the polarizabilities of nihonium (Nh, $Z=113$), computed with the FSCC method and a custom Fægri basis set.9 Dzuba et al. subsequently computed the value for Nh using the SD+CI method in a Dirac framework with Breit interaction and high-order QED correlation, agreeing with the previous work.10 Fleig and Sadlej investigated elements fluorine (F, $Z=9$), bromine (Br, $Z=35$), iodine (I, $Z=53$), and astatine (At, $Z=85$) with two-component spin-orbit coupling CI calculations,11 while tennessine (Ts, $Z=117$) currently only has an empirically estimated value.12 No prior study has provided information on the contribution of the SOC effect to Groups 15–16 atoms, with the exception of moscovium (Mc, $Z=115$), which was examined using the SD+CI approach by Dzuba et al.10 Moreover, errors associated with the recommended values for Groups 15–16 atoms are still significant.3 Currently, no theoretical value exists for Lv.
@@
The main goal of this study is to provide a comprehensive and consistent dataset of dipole polarizabilities for main-group elements. To achieve this, I employed relativistic single-reference coupled cluster (CC) methods and MRCI methods within the Dirac-Coulomb Hamiltonian framework. I included more correlated electrons and virtual orbitals/spinors in correlation energy calculations, using larger basis sets than previous theoretical studies. Additionally, I investigated the impact of electronic correlation and relativistic effects on dipole polarizabilities through a quantitative approach. It should be noted that the data for light atoms ($Z<40$) are also computed in this work. However, their values have been extensively investigated, and improving their theoretical values is beyond the scope of this study. Readers who are interested in these data can refer to Ref. 3 and its latest version.13
@@ 2.1 Relativistic framework
In this study, all energies were calculated within a fully relativistic framework using a four-component relativistic formalism. The relativistic effects can be understood as a combination of SOC effects and scalar-relativistic effects, which account for the contraction or decontraction of the radial electron density.14
@@ 2.1.1 Four-component Dirac-Coulomb Hamiltonian
An accurate standard for four-component relativistic electronic-structure theory is based on the Dirac-Coulomb-Breit (DCB) Hamiltonian:15
@@
In the context of two-electron interactions, relativistic corrections up to the second order, specifically the Coulomb-Breit interaction in the Coulomb gauge, are represented by the following equation:16
@@ 2.1.2 Exact Two-Component Dirac Hamiltonian
The high computational costs associated with four-component relativistic calculations have prompted the development of methods that use two-component relativistic Hamiltonians. Examples of these methods include the second-order Douglas-Kroll-Hess (DKH) Hamiltonian 17, 18, 19 and zeroth-order regular approximation (ZORA) Hamiltonians.20, 21, 22 However, these methods are all limited to finite orders. Ilias et al. proposed an infinite-order two-component (IOTC) relativistic Hamiltonian, also known as the exact two-component (X2C) method.23 In the X2C method, the relativistic effects can be split into a spin-free part and a spin-dependent part, which correspond to scalar-relativistic and the SOC effects, respectively. In this work, the spin-free X2C method is employed to calculate both non-relativistic and scalar-relativistic properties. More details about the X2C method can be found in recent reviews.16, 24
@@ 2.2 Relativistic and Correlation Effects
The Dirac Hartree-Fock (DHF) method was utilized as the uncorrelated method. In this study, the terms “orbital” and “spinor” are used to describe the quantum state of an electron without and with considering SOC effects. Based on these reference states, correlated energies were calculated. Examples of these methods include the second-order Møller-Plesset (MP2) perturbation theory,25 coupled-cluster models CCSD and CCSD(T), where SD stands for single and double excitations, and “(T)” represents perturbed triple excitations,26 and the MRCI method with single and double substitutions (MRCISD).27, 28, 29, 30
@@
In addition, the SOC effect on dipole polarizabilities can be then defined as:8, 14
@@ 2.3 Open-shell methods
For the spin-dependent coupling case ($jj$ representation), the averaged polarizability $\bar{\alpha}^{J}$ for the ground state is determined by:14
@@
where $\alpha(J,M_{J})$ denotes the polarizabilities for each $J,M_{J}$ component. However, for spin-free cases ($LS$ representation), the equation is analogous to Eq. (10) with $J$ replaced by $L$:14
@@
In this work, only $p$-block open-shell elements are studied. The atomic $p$ spinors in the relativistic case are linear combinations of the real $p$ orbitals in the non-relativistic case:11
@@ 2.3.1 Group 13 elements
The spin-orbit components of Group 13 atomic states can be represented as the $p^{1}$ states using atomic spinors $\phi(j,m_{j})$ due to the presence of only one electron in $p$-spinor. In the ground state, only the spinors $\phi\left(\frac{1}{2},\pm\frac{1}{2}\right)$ are occupied. When an external perturbation, such as an electric field along the $z$ axis, is applied, the two degenerate spinors $\phi\left(\frac{3}{2},\pm\frac{3}{2}\right)$ and spinors $\phi\left(\frac{3}{2},\pm\frac{1}{2}\right)$ split, resulting in a non-degenerate state. The energy shift is dependent on the $p_{z}$-component and generally, spinors with a larger $p_{z}$ contribution will experience a larger energy shift than those with a smaller contribution.11 However, as weak electric fields are typically applied, the energy shift is much smaller in absolute value than the spin-orbit splitting. As a result, the $J=\frac{3}{2},M_{J}=\pm\frac{3}{2}$ states will have a smaller dipole polarizability than the $J=\frac{3}{2},M_{J}=\pm\frac{1}{2}$ states.
Based on the analysis in Ref. 8, the value $\bar{\alpha}_{J=3/2}-\alpha_{J=1/2}$ can be interpreted as the effect of both SOC and the contraction (decontraction) of the $p_{1/2}$ ($p_{3/2}$) spinor in total. The SOC effect can also be determined by comparing the $M_{L}$-averaged polarizability, calculated using SR-CC methods, to the value of the $J=\frac{1}{2}$ ground state computed using the DC-CC method. Moreover, the DC-CI method allows for the computation of excited-state polarizabilities, facilitating the study of spin-orbit resolved polarizabilities for these excited state components.
@@ 2.3.3 Group 15 elements
The DC-CI calculations were performed to investigate the electron-correlation and the SOC effect on polarizabilities. Additionally, the polarizabilities of different $M_{J}$ sub-states were studied. In practice, DC-CI eigenvectors are often represented by string occupation and the spin-projection value $M_{K}$, which is an integer multiple of $\frac{1}{2}$ for a given determinant.11 As a result, the different components of ${{}^{4}}P_{3/2}$ are not pure states, and may be combined with high $M_{K}$ states that have the same $M_{J}$. To address this, the method proposed by Fleig et al. 11 was used to approximately specify the $M_{J}$ state. However, since the primitive determinant is determined by $M_{J}=\pm 3/2$ and $M_{J}=\pm 1/2$ for $M_{K}=\frac{1}{2}$, the polarizabilities of these states are thus approximately considered to represent ground-state values. In this work, the states $J=\frac{3}{2},M_{J}=\pm\frac{1}{2}$ refer to the states with lower energy compared to the states $J=\frac{3}{2},M_{J}=\pm\frac{3}{2}$ when an external electric field is applied along the $z$-axis.
@@ 2.3.5 Group 17 elements
The spin-orbit components of the ground states of halogen atoms can be represented as the inverse of the $p^{5}$ hole states by using atomic spinors $\phi(j,m_{j})$ as there is only one hole in the $p$-spinor for these atoms.11 The polarizabilities of different components of the ${{}^{2}}P_{3/2}$ states were determined using a similar procedure to that used for Group 15 elements. The DC-CI method was used to calculate the polarizabilities for states $M_{J}=\pm 3/2$ and $M_{J}=\pm 1/2$, and the $M_{J}$-averaged polarizability is approximated as the ground-state value. Moreover, the computational values for states ${{}^{2}}P_{1/2}$ are also computed.
@@ 2.4 Finite-field methods
The static dipole polarizabilities presented in this work were computed using finite-field methods. 31. The energy of an atom in a homogeneous electric field of strength $F_{z}$ along the $z$-axis is represented as follows:31
@@
The least-squares procedure is used instead of numerical differentiation for reasons discussed in Ref. 32. In practice, this procedure could lead to an unreasonable $\gamma$, including extremely positive or negative values. In these cases, an expansion only including $\alpha$ is also used:
@@ 2.5 Uncertainty estimation
The uncertainty estimation is analyzed empirically based on the composite scheme, which has been widely employed in previous studies to obtain the final value of an atomic property, denoted as $P_{\text{final}}$.33, 14, 34 It should be noted that the $P_{\text{final}}$ values are not provided in this work due to their empirical nature and are only used for uncertainty estimation empirically.
@@ 2.5.1 Composite scheme of the CC method
where $P_{\text{CCSD}}$, $P_{\text{CCSD(T)}}$, $P_{\text{CCSDT}}$, and $P_{\text{CCSDTQ}}$ are the values calculated for the property using CCSD, CCSD(T), CCSDT (CC singles, doubles, and triples), and CCSDTQ (CC singles, doubles, triples, and quadruples), respectively. $\Delta P_{\text{basis}}$ is the error due to the use of the finite-size basis functions for computing $P_{\text{CCSD}}$ or $P_{\text{CCSD(T)}}$.35, 34 $\Delta P_{\text{core}}$ is the correction due to the more outer-core shells, while $\Delta P_{\text{vir}}$ is the possible error due to the neglected contributions from the frozen high-lying virtual orbitals/spinors.35 $P_{\text{CCSD}}$ or $P_{\text{CCSD(T)}}$ is the value for property $P$ using the most accurate CCSD or CCSD(T) method in this work. $\Delta P_{\text{fitting}}$ is the numerical error resulting from the fitting procedure using Eq. (14) or Eq. (15). For completeness, $\Delta P_{\text{others}}$ corresponds to all other corrections, e.g., the difference between $\Delta P_{\text{(T)}}$ and $\Delta P_{\text{T}}$ when $\Delta P_{\text{(T)}}$ is used, or other relativistic effects, including Gaunt, Breit, or QED interactions.
@@ 2.5.2 Composite scheme of the MRCI method
where $P_{\text{MRCISD}}$, $P_{\text{MRCISDT}}$, and $P_{\text{MRCISDTQ}}$ are the values calculated for the property using MRCISD, MRCISDT (MRCI singles, doubles, and triples), and MRCISDTQ (MRCI singles, doubles, triples, and quadruples), respectively. $\Delta P_{\text{basis}}$ is the error due to the use of the finite-size basis functions for computing $P_{\text{MRCISD}}$.35, 34 $\Delta P_{\text{core}}$ is the correction due to the more outer-core shells, while $\Delta P_{\text{vir}}$ is the possible error due to the neglected contributions from the frozen high-lying virtual spinors.35 $\Delta P_{\text{fitting}}$ is the numerical error resulting from the fitting procedure using Eq. (14) or Eq. (15). Similarly, $\Delta P_{\text{others}}$ corresponds to all other corrections as in CC calculations.

### v2

@@ 1 Introduction
The electric dipole polarizability quantifies the response of a system’s electron density to an external electric field, leading to induced dipoles. This property plays a critical role in many applications, such as determining atomic scattering cross sections, refractive indices, dielectric constants, interatomic interactions, and the development of polarizable force fields used in molecular simulations.1 Moreover, an accurate value of the dipole polarizability helps benchmark other methods, such as density-functional theory (DFT).2
The accurate determination of polarizabilities is essential not only for atomic and molecular physics but also for precision technologies. For example, in atomic clocks, the leading-order Stark shift and black-body radiation (BBR) shift, which are critical for timekeeping accuracy, depend on precise values of atomic dipole polarizability.3, 4 The BBR shift has contributed to the primary bottleneck of overall uncertainty in strontium (Sr, $Z=38$) and ytterbium (Yb, $Z=70$) clocks for several years at the $10^{-16}$ level.5, 6, 7 To further improve the accuracy of atomic clocks, higher precision in the differential static polarizability of the clock states is required.7
@@
The determination of accurate atomic polarizabilities requires high-level electron correlation treatment within a relativistic framework, particularly for heavy atoms. Over the past two decades, researchers have advanced techniques to incorporate spin-orbit coupling (SOC) effects in dipole polarizability calculations. For closed-shell atoms and open-shell atoms with only one open-shell electron, the coupled cluster theory with single and double substitutions (CCSD) with perturbative triples, CCSD(T), is usually employed within a fully relativistic Dirac-Coulomb framework, where both scalar-relativistic effects and SOC effects can be included. Due to the SOC effect, atoms in Group 14 are also treated as closed-shell atoms. However, for open-shell atoms with more than one open-shell electron in Groups 15–17, the coupled cluster theory is not yet well-developed at the fully relativistic level, and an alternative is multi-reference configuration interaction (MRCI).1
@@
Previous studies have thoroughly investigated the polarizabilities considering SOC effects at the CCSD or CCSD(T) level for cesium (Cs, $Z=55$)10, 11 and francium (Fr, $Z=87$)10, 12, 13 in Group 1. Fleig et al. examined the spin-orbit-resolved static polarizabilities of Group 13 atoms, including gallium (Ga, $Z=31$), indium (In, $Z=49$), and thallium (Tl, $Z=81$), using four-component configuration interaction (CI) and coupled cluster methods with ANO-RCC basis sets in the Dirac-Coulomb framework.14 Borschevsky et al. later improved these results using the Fock space coupled cluster (FSCC) with more extensive Fægri basis sets.10 Pershina et al. first reported the polarizabilities of nihonium (Nh, $Z=113$), computed with the FSCC method and a custom Fægri basis set.15 Dzuba et al. subsequently computed the value for Nh using the SD+CI method in a Dirac framework with Breit interaction and high-order quantum electrodynamics (QED) correlation, agreeing with the previous work.16 Fleig and Sadlej investigated elements fluorine (F, $Z=9$), bromine (Br, $Z=35$), iodine (I, $Z=53$), and astatine (At, $Z=85$) with two-component spin-orbit coupling CI calculations,17 while tennessine (Ts, $Z=117$) currently only has an empirically estimated value.18 No prior study has provided information on the contribution of the SOC effect to Groups 15–16 atoms, with the exception of moscovium (Mc, $Z=115$), which was examined using the SD+CI approach by Dzuba et al.16 Moreover, errors associated with the recommended values for Groups 15–16 atoms are still significant.1 Currently, no theoretical value exists for Lv.
@@
The main goal of this study is to provide a comprehensive and consistent dataset of dipole polarizabilities for main-group elements. To achieve this, I employed relativistic single-reference coupled cluster (CC) methods and MRCI methods within the Dirac-Coulomb Hamiltonian framework. I included more correlated electrons and virtual orbitals/spinors in correlation energy calculations, using larger basis sets than previous theoretical studies. Additionally, I investigated the impact of electronic correlation and relativistic effects on dipole polarizabilities through a quantitative approach. It should be noted that the data for light atoms ($Z<40$) are also computed in this work. However, their values have been extensively investigated, and improving their theoretical values is beyond the scope of this study. Readers who are interested in these data can refer to Ref. 1 and its latest version.19
@@ 2.1 Relativistic framework
In this study, all energies were calculated within a fully relativistic framework using a four-component relativistic formalism. The relativistic effects can be understood as a combination of SOC effects and scalar-relativistic effects, which account for the contraction or decontraction of the radial electron density.20
@@ 2.1.1 Four-component Dirac-Coulomb Hamiltonian
An accurate standard for four-component relativistic electronic-structure theory is based on the Dirac-Coulomb-Breit (DCB) Hamiltonian:21
@@
In the context of two-electron interactions, relativistic corrections up to the second order, specifically the Coulomb-Breit interaction in the Coulomb gauge, are represented by the following equation:22
@@ 2.1.2 Exact Two-Component Dirac Hamiltonian
The high computational costs associated with four-component relativistic calculations have prompted the development of methods that use two-component relativistic Hamiltonians. Examples of these methods include the second-order Douglas-Kroll-Hess (DKH) Hamiltonian 23, 24, 25 and zeroth-order regular approximation (ZORA) Hamiltonians.26, 27, 28 However, these methods are all limited to finite orders. Ilias et al. proposed an infinite-order two-component (IOTC) relativistic Hamiltonian, also known as the exact two-component (X2C) method.29 In the X2C method, the relativistic effects can be split into a spin-free part and a spin-dependent part, which correspond to scalar-relativistic and the SOC effects, respectively. In this work, the spin-free X2C method is employed to calculate both non-relativistic and scalar-relativistic properties. More details about the X2C method can be found in recent reviews.22, 30
@@ 2.2 Relativistic and Correlation Effects
The Dirac Hartree-Fock (DHF) method was utilized as the uncorrelated method. In this study, the terms “orbital” and “spinor” are used to describe the quantum state of an electron without and with considering SOC effects. Based on these reference states, correlated energies were calculated. Examples of these methods include the second-order Møller-Plesset (MP2) perturbation theory,31 coupled-cluster models CCSD and CCSD(T), where SD stands for single and double excitations, and “(T)” represents perturbed triple excitations,32 and the MRCI method with single and double substitutions (MRCISD).33, 34, 35, 36
@@
In addition, the SOC effect on dipole polarizabilities can be then defined as:14, 20
@@ 2.3 Open-shell methods
For the spin-dependent coupling case ($jj$ representation), the averaged polarizability $\bar{\alpha}^{J}$ for the ground state is determined by:20
@@
where $\alpha(J,M_{J})$ denotes the polarizabilities for each $J,M_{J}$ component. However, for spin-free cases ($LS$ representation), the equation is analogous to Eq. (10) with $J$ replaced by $L$:20
@@
In this work, only $p$-block open-shell elements are studied. The atomic $p$ spinors in the relativistic case are linear combinations of the real $p$ orbitals in the non-relativistic case:17
@@ 2.3.1 Group 13 elements
The spin-orbit components of Group 13 atomic states can be represented as the $p^{1}$ states using atomic spinors $\phi(j,m_{j})$ due to the presence of only one electron in $p$-spinor. In the ground state, only the spinors $\phi\left(\frac{1}{2},\pm\frac{1}{2}\right)$ are occupied. When an external perturbation, such as an electric field along the $z$ axis, is applied, the two degenerate spinors $\phi\left(\frac{3}{2},\pm\frac{3}{2}\right)$ and spinors $\phi\left(\frac{3}{2},\pm\frac{1}{2}\right)$ split, resulting in a non-degenerate state. The energy shift is dependent on the $p_{z}$-component and generally, spinors with a larger $p_{z}$ contribution will experience a larger energy shift than those with a smaller contribution.17 However, as weak electric fields are typically applied, the energy shift is much smaller in absolute value than the spin-orbit splitting. As a result, the $J=\frac{3}{2},M_{J}=\pm\frac{3}{2}$ states will have a smaller dipole polarizability than the $J=\frac{3}{2},M_{J}=\pm\frac{1}{2}$ states.
Based on the analysis in Ref. 14, the value $\bar{\alpha}_{J=3/2}-\alpha_{J=1/2}$ can be interpreted as the effect of both SOC and the contraction (decontraction) of the $p_{1/2}$ ($p_{3/2}$) spinor in total. The SOC effect can also be determined by comparing the $M_{L}$-averaged polarizability, calculated using SR-CC methods, to the value of the $J=\frac{1}{2}$ ground state computed using the DC-CC method. Moreover, the DC-CI method allows for the computation of excited-state polarizabilities, facilitating the study of spin-orbit resolved polarizabilities for these excited state components.
@@ 2.3.3 Group 15 elements
The DC-CI calculations were performed to investigate the electron-correlation and the SOC effect on polarizabilities. Additionally, the polarizabilities of different $M_{J}$ sub-states were studied. In practice, DC-CI eigenvectors are often represented by string occupation and the spin-projection value $M_{K}$, which is an integer multiple of $\frac{1}{2}$ for a given determinant.17 As a result, the different components of ${{}^{4}}P_{3/2}$ are not pure states, and may be combined with high $M_{K}$ states that have the same $M_{J}$. To address this, the method proposed by Fleig et al. 17 was used to approximately specify the $M_{J}$ state. However, since the primitive determinant is determined by $M_{J}=\pm 3/2$ and $M_{J}=\pm 1/2$ for $M_{K}=\frac{1}{2}$, the polarizabilities of these states are thus approximately considered to represent ground-state values. In this work, the states $J=\frac{3}{2},M_{J}=\pm\frac{1}{2}$ refer to the states with lower energy compared to the states $J=\frac{3}{2},M_{J}=\pm\frac{3}{2}$ when an external electric field is applied along the $z$-axis.
@@ 2.3.5 Group 17 elements
The spin-orbit components of the ground states of halogen atoms can be represented as the inverse of the $p^{5}$ hole states by using atomic spinors $\phi(j,m_{j})$ as there is only one hole in the $p$-spinor for these atoms.17 The polarizabilities of different components of the ${{}^{2}}P_{3/2}$ states were determined using a similar procedure to that used for Group 15 elements. The DC-CI method was used to calculate the polarizabilities for states $M_{J}=\pm 3/2$ and $M_{J}=\pm 1/2$, and the $M_{J}$-averaged polarizability is approximated as the ground-state value. Moreover, the computational values for states ${{}^{2}}P_{1/2}$ are also computed.
@@ 2.4 Finite-field methods
The static dipole polarizabilities presented in this work were computed using finite-field methods. 37. The energy of an atom in a homogeneous electric field of strength $F_{z}$ along the $z$-axis is represented as follows:37
@@
The least-squares procedure is used instead of numerical differentiation for reasons discussed in Ref. 38. In practice, this procedure could lead to an unreasonable $\gamma$, including extremely positive or negative values. In these cases, an expansion only including $\alpha$ is also used:
@@ 2.5 Uncertainty estimation
The uncertainty estimation is analyzed empirically based on the composite scheme, which has been widely employed in previous studies to obtain the final value of an atomic property, denoted as $P_{\text{final}}$.39, 20, 40 It should be noted that the $P_{\text{final}}$ values are not provided in this work due to their empirical nature and are only used for uncertainty estimation empirically.
@@ 2.5.1 Composite scheme of the CC method
where $P_{\text{CCSD}}$, $P_{\text{CCSD(T)}}$, $P_{\text{CCSDT}}$, and $P_{\text{CCSDTQ}}$ are the values calculated for the property using CCSD, CCSD(T), CCSDT (CC singles, doubles, and triples), and CCSDTQ (CC singles, doubles, triples, and quadruples), respectively. $\Delta P_{\text{basis}}$ is the error due to the use of the finite-size basis functions for computing $P_{\text{CCSD}}$ or $P_{\text{CCSD(T)}}$.4, 40 $\Delta P_{\text{core}}$ is the correction due to the more outer-core shells, while $\Delta P_{\text{vir}}$ is the possible error due to the neglected contributions from the frozen high-lying virtual orbitals/spinors.4 $P_{\text{CCSD}}$ or $P_{\text{CCSD(T)}}$ is the value for property $P$ using the most accurate CCSD or CCSD(T) method in this work. $\Delta P_{\text{fitting}}$ is the numerical error resulting from the fitting procedure using Eq. (14) or Eq. (15). For completeness, $\Delta P_{\text{others}}$ corresponds to all other corrections, e.g., the difference between $\Delta P_{\text{(T)}}$ and $\Delta P_{\text{T}}$ when $\Delta P_{\text{(T)}}$ is used, or other relativistic effects, including Gaunt, Breit, or QED interactions.
@@ 2.5.2 Composite scheme of the MRCI method
where $P_{\text{MRCISD}}$, $P_{\text{MRCISDT}}$, and $P_{\text{MRCISDTQ}}$ are the values calculated for the property using MRCISD, MRCISDT (MRCI singles, doubles, and triples), and MRCISDTQ (MRCI singles, doubles, triples, and quadruples), respectively. $\Delta P_{\text{basis}}$ is the error due to the use of the finite-size basis functions for computing $P_{\text{MRCISD}}$.4, 40 $\Delta P_{\text{core}}$ is the correction due to the more outer-core shells, while $\Delta P_{\text{vir}}$ is the possible error due to the neglected contributions from the frozen high-lying virtual spinors.4 $\Delta P_{\text{fitting}}$ is the numerical error resulting from the fitting procedure using Eq. (14) or Eq. (15). Similarly, $\Delta P_{\text{others}}$ corresponds to all other corrections as in CC calculations.